OpenStudy (anonymous):

Angle BAD is bisected by ray AC. Find the measures of Angle BAC and Angle CAD if the measure of Angle BAC is 11x + 5 and the measure of Angle CAD is 7x + 8. x = Measure of Angle BAC = o Measure of Angle CAD = o

OpenStudy (anas.p):

Remember that bisector means dividing in equal halves... what does that tell us about angle BAC and angle CAD?

OpenStudy (anas.p):

 look carefully and tell me about the relation between BAC and CAD

OpenStudy (anas.p):

CAD=BAC

OpenStudy (anas.p):

because they are equal halves

OpenStudy (anas.p):

no... Equate the two equations... 11x+5=7x+8

OpenStudy (anonymous):

Is it 11x+5+7x+8=180 or 11x+5+7x+8=90

OpenStudy (anas.p):

get it? you equate the two equations...